OpenStudy (anonymous):

Solve x^2 + 6x - 5 = 0 by completing the square. Which equation is used in the process?

OpenStudy (anonymous):

add 5 to both sides and get \[x^2+6x=5\] then take half of 6 namely 3, square it and get 9 and write \[(x+3)^2=5+9\] or \[(x+3)^2=14\] take the square root of both sides, don't forget the \(\pm\) and get \[x+3=\pm\sqrt{14}\] and finally solve for \(x\) to get \[x=-3\pm\sqrt{14}\]